# Mixture of Experts Meets Prompt-Based Continual Learning

### v1

@@ 1 Introduction
Humans possess a remarkable ability to learn continuously by integrating new skills and knowledge while retaining past experiences. However, current AI models often fail to retain this ability. Unlike humans, they often suffer from catastrophic forgetting [24, 25, 27], a phenomenon where they struggle to retain knowledge from previous tasks while learning new ones. Inspired by human learning, Continual Learning [2, 24, 23, 1] is an ongoing field that aims to train a model across a sequence of tasks while mitigating this challenge. Traditional continual learning methods often rely on storing past data for fine-tuning, which can raise concerns about memory usage and privacy [5, 32, 42]. To address these limitations, prompt-based approaches have emerged as a promising alternative within rehearsal-free continual learning. By attaching prompts - small sets of learnable parameters - to a frozen pre-trained model, these approaches enable efficient adaptation to new tasks with minimal modifications to the underlying model [47, 21, 50]. The effectiveness of prompt-based methods has been demonstrated by several recent works achieving state-of-the-art performance on various continual learning benchmarks [40, 44, 45].
@@
Our contributions can be summarized as follows: (1) We reveal a novel connection between self-attention and a mixture of experts, providing a fresh perspective on prompt-based continual learning approaches; (2) Leveraging this insight, we propose Non-linear Residual Gates (NoRGa), an innovative gating mechanism that enhances continual learning performance while maintaining parameter efficiency, and provide a theoretical justification for this improvement; (3) Extensive experiments across various continual learning benchmarks and pretraining settings demonstrate that our approach achieves state-of-the-art performance compared to existing methods.
@@ 2 Background and Related Works
Prompt-based continual learning. Prompt-based approaches have emerged as a promising alternative within rehearsal-free continual learning [50, 43]. In vision tasks, prompt-based methods often leverage a pre-trained ViT as a feature extractor $f_{\theta}$, with its parameters $\theta$ typically frozen. These methods enhance the model by introducing prompts, small sets of learnable parameters that influence the operations of the MSA layer [44]. Prompts are strategically injected into the query, key, and value matrices to guide the ViT in learning new tasks. We denote the prompt parameters by $\bm{p}\in\mathbb{R}^{L_{p}\times d}$, where $L_{p}$ is the sequence length and $d$ is the embedding dimension. Previous work [44] outlines two main prompt-based approaches: Prompt Tuning (ProT) [20] and Prefix Tuning (PreT) [21]. While Prompt Tuning directly concatenates the same prompt parameter $\bm{p}$ to the query, key, and value, Prefix Tuning divides $\bm{p}$ into prefixes $\{\bm{p}^{K},\bm{p}^{V}\}\in\mathbb{R}^{\frac{L_{p}}{2}\times d}$ and appends it to the key and value vectors:
@@
In this study, we focus on Prefix Tuning as our primary prompt-based methodology and follow the framework presented in HiDe-Prompt [40]. During training, HiDe-Prompt co-optimizes task-specific prompts $\bm{p}_{t}$ and model’s output layer parameters $\psi$ for each new task $t$ using the WTP objective. These prompts are stored within a prompt pool $\mathbf{P}=\{\bm{p}_{1},...,\bm{p}_{T}\}$. At test time, a separate lightweight auxiliary output layer $\hat{h}_{\omega}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{T}$, trained with the TII objective, takes the uninstructed representation $f_{\theta}(x)$ of a new data point $\bm{x}$ as input to infer the task identity, guiding the selection of the most suitable prompt $\bm{p}_{k}$ from the prompt pool $\mathbf{P}$. Subsequently, the final prediction is given as $\hat{y}=h_{\psi}(f_{\theta}({\bm{x}},\bm{p}_{k}))$. For further details, please refer to Appendix C.
@@
where $G({\bm{h}}):=\mathrm{softmax}(s_{1}({\bm{h}}),\ldots,s_{N}(h))$. Building on this concept, works by [10, 34] established the MoE layer as a fundamental building block to scale up model capacity efficiently.
@@ 3.1 Mixture of Experts Meets Attention
for $i\in[N]$. Expanding on equation (8), we can discern that each attention head within the MSA layer implicitly embodies a special mixture of experts architecture. This architecture encompasses $N$ MoE models, each featuring its own quadratic gating function $G_{i}$. However, instead of employing $N^{2}$ separate expert networks for each model, this architecture utilizes $N$ shared linear expert networks $f_{j}$ for $j\in[N]$, significantly reducing the number of parameters. Notably, each expert network and its corresponding gating function process the entire input sequence directly, rather than individual embedding ${\bm{x}}_{i}$ as in traditional MoE layers [34]. This connection between self-attention and mixture of experts motivates us to explore how prompt-based techniques can be viewed through this lens.
@@ 3.2 Prefix Tuning via the Perspective of Mixture of Experts
Building on the connection between self-attention and mixture of experts, we propose that applying prefix tuning can be interpreted as the introduction of new experts to customize the pre-trained model for a specific task. Specifically, similar to Section 3.1, we consider the $l$-th head within the MSA layer. We denote $\bm{p}^{K}=\left[\bm{p}_{1}^{K},\dots,\bm{p}_{L}^{K}\right]^{\top}\in\mathbb{R}^{L\times d}$, $\bm{p}^{V}=\left[\bm{p}_{1}^{V},\dots,\bm{p}_{L}^{V}\right]^{\top}\in\mathbb{R}^{L\times d}$, where $L=\frac{L_{p}}{2}$. We define new prefix experts $f_{N+j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$ along with their corresponding new score functions $s_{i,N+j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}$ as follows:
@@
In the context of continual learning, the pre-trained experts serve as a knowledge base, while Prefix Tuning augments it with task-specific knowledge encoded in new experts. Moreover, we draw a parallel between the pre-trained experts and the G(eneral)-Prompt utilized in DualPrompt, which captures task-agnostic information [44]. Both are shared across tasks, making them useful for prediction, especially when task identification is incorrect. Notably, the new experts achieve their efficiency through simple linear gating functions and independence from the input, unlike the pre-trained experts. For simplicity, we call the MoE model (11) as linear gating prefix MoE.
@@ 4 Non-linear Residual Gate Meets Prefix Tuning
As discussed earlier, prefix tuning introduces additional experts within MoE framework, resulting in the linear gating prefix MoE model. However, as outlined in Appendix A, this approach suffers from suboptimal sample efficiency for parameter estimation. To overcome this and enhance overall CIL performance, we propose an innovative approach that significantly improves sample efficiency while promoting WTP performance in Section 4.1 and provide theoretical explanations in Section 4.2.
@@ 4.1 NoRGa: Non-linear Residual Gate
where $\alpha,\tau\in\mathbb{R}$ are scalar factors, and $\sigma$ is a non-linear activation function. The new score function in equation (12) consists of a linear and a non-linear component. We call the new prefix MoE model with score functions (12) as non-linear residual gating prefix MoE.
The use of a non-linear activation function here is motivated by the algebraic independence condition in Definition 4.2 to theoretically guarantee the optimal sample efficiency of expert and parameter estimations (cf. Theorem 4.3). It’s worth noting that removing the linear component $s_{i,N+j}({\bm{X}})$ in the score function (12) could potentially lead to the vanishing gradient problem during training. To mitigate this challenge, we incorporate a residual connection [13] into the formulation. Our modification introduces minimal additional parameters ($\alpha$ and $\tau$) compared to the original score function, ensuring parameter efficiency. This is particularly crucial in continual learning scenarios where the number of parameters grows with each new task. Despite its simplicity, our modification can significantly enhance sample efficiency and promote more reasonable parameter estimation, as demonstrated in our theoretical analysis in Section 4.2. Within the HiDe-Prompt framework, task-specific prompt parameters are trained using the WTP objective for each new task. Consequently, our modification leads to better parameter estimation, which directly contributes to improved WTP performance, ultimately improving overall continual learning efficacy. Here, we evaluated $\sigma$ with $\tanh$, $\mathrm{sigmoid}$, and $\mathrm{GELU}$, finding $\tanh$ to perform well in most cases.
@@ 5 Experiments
Datasets We evaluate various continual learning methods on widely used CIL benchmarks, including Split CIFAR-100 [18] and Split ImageNet-R [18], consistent with prior work [40]. We further explore the model’s performance on fine-grained classification tasks with Split CUB-200 [39] and large inter-task differences with 5-Datasets [9]. Please refer to Appendix D for more details.
Evaluation Metrics We utilize several established metrics described in [41]. These include: final average accuracy (FA), which represents the average accuracy after the final task; cumulative average accuracy (CA), which refers to the historical average accuracy; and average forgetting measure (FM). We give more emphasis to FA and CA due to their comprehensiveness, as noted in [35].
Baselines For the comparison, we select representative prompt-based approaches including L2P [45], DualPrompt [44], CODA-Prompt [35], S-Prompt [43], and HiDe-Prompt [40]. Additionally, in line with [40], we utilize the checkpoints of ViT that use supervised pretraining of Imagenet-21K (denoted as Sup-21K), and some self-supervised pretraining such as iBOT-21K, iBOT-1K [51], DINO-1K [4], and MoCo-1K [7]. For implementation details, see Appendix D.
@@
Main Results. In Table 1, we evaluate the performance of various continual learning methods on the Split CIFAR-100 and Split ImageNet-R datasets using diverse pre-trained models. NoRGa achieves state-of-the-art FA and CA across all datasets and models, consistently outperforming HiDe-Prompt. On Sup-21K, NoRGa shows impressive FA results on CIFAR-100 and ImageNet-R. It also maintains the highest CA, with significant gaps of 1.80% and 2.92% on CIFAR-100 and ImageNet-R, respectively, compared to HiDe-Prompt. These results highlight NoRGa’s ability to retain knowledge and sustain high accuracy throughout the continual learning process. Additionally, NoRGa exhibits minimal forgetting behavior, as evidenced by the low FM values on both datasets. When evaluated with self-supervised pre-training models, NoRGa continues to excel, outperforming HiDe-Prompt by up to 1.95% and 3.66% on the two datasets in terms of FA. We further investigate two scenarios: fine-grained classification tasks and large inter-task differences through experiments on Split CUB-200 and 5-Datasets, respectively, in Table 2. On Split CUB-200, NoRGa achieves an impressive FA with a gap of 4.34% with Sup-21k and 2.46% with iBOT-21k compared to HiDe-Prompt. Similarly, on the 5-Datasets, NoRGa maintains its superiority with the highest FA. These results underscore NoRGa’s robustness and effectiveness across diverse datasets.
Ablation Study. To assess the impact of non-linear activation functions on NoRGa’s performance, we evaluated the model’s behavior with different choices for the activation function $\sigma$, including $\tanh$, $\mathrm{sigmoid}$, and $\mathrm{GELU}$ in Table 3. The results show that NoRGa achieves state-of-the-art performance on both Split CIFAR-100 and Split CUB-200 datasets with all three activation functions. These findings suggest that NoRGa exhibits robustness to the choice of non-linear activation within a reasonable range. While all functions perform well, the $\tanh$ activation function demonstrates generally strong performance across scenarios. To further validate NoRGa’s effectiveness in improving WTP performance, we perform experiments on the task-incremental learning setting (see Appendix E).
@@ 6 Conclusion
In this supplementary material, we first investigate the statistical suboptimality of the Linear Gating Prefix MoE Model (11). Next, we provide proofs for the theoretical results presented in Section 4.2. Subsequently, Appendix D specifies the details for the experiments conducted in Section 5. Finally, Appendix E presents further experiments on the task-incremental learning setting to empirically demonstrate the benefits of using our proposed Non-linear Residual Gating Prefix MoE (12) over the Linear Gating Prefix MoE Model.

### v2

@@ 1 Introduction
Humans possess a remarkable ability to learn continuously by integrating new skills and knowledge while retaining past experiences. However, current AI models often fail to retain this ability. Unlike humans, they often suffer from catastrophic forgetting [28, 30, 32, 38], a phenomenon where they struggle to retain knowledge from previous tasks while learning new ones. Inspired by human learning, Continual Learning [2, 28, 43, 1, 12] is an ongoing field that aims to train a model across a sequence of tasks while mitigating this challenge. Traditional continual learning methods often rely on storing past data for fine-tuning, which can raise concerns about memory usage and privacy [5, 39, 51]. To address these limitations, prompt-based approaches have emerged as a promising alternative within rehearsal-free continual learning. By attaching prompts - small sets of learnable parameters - to a frozen pre-trained model, these approaches enable efficient adaptation to new tasks with minimal modifications to the underlying model [56, 26, 61]. The effectiveness of prompt-based methods has been demonstrated by several recent works achieving state-of-the-art performance on various continual learning benchmarks [49, 53, 54].
@@
Our contributions can be summarized as follows: (1) We reveal a novel connection between self-attention and a mixture of experts, providing a fresh perspective on prompt-based continual learning approaches; (2) Leveraging this insight, we propose Non-linear Residual Gates (NoRGa), an innovative gating mechanism that enhances continual learning performance while maintaining parameter efficiency, and provide a theoretical justification for this improvement; (3) Extensive experiments across various continual learning benchmarks and pre-training settings demonstrate that our approach achieves state-of-the-art performance compared to existing methods.
@@ 2 Background and Related Works
Prompt-based continual learning. Prompt-based approaches have emerged as a promising alternative within rehearsal-free continual learning [61, 52, 44]. In vision tasks, prompt-based methods often leverage a pre-trained ViT as a feature extractor $f_{\theta}$, with its parameters $\theta$ typically frozen. These methods enhance the model by introducing prompts, small sets of learnable parameters that influence the operations of the MSA layer [53]. Prompts are strategically injected into the query, key, and value matrices to guide the ViT in learning new tasks. We denote the prompt parameters by $\bm{p}\in\mathbb{R}^{L_{p}\times d}$, where $L_{p}$ is the sequence length and $d$ is the embedding dimension. Previous work [53] outlines two main prompt-based approaches: Prompt Tuning (ProT) [25] and Prefix Tuning (PreT) [26]. While Prompt Tuning directly concatenates the same prompt parameter $\bm{p}$ to the query, key, and value, prefix tuning divides $\bm{p}$ into prefixes $\{\bm{p}^{K},\bm{p}^{V}\}\in\mathbb{R}^{\frac{L_{p}}{2}\times d}$ and appends it to the key and value vectors:
@@
In this study, we focus on prefix tuning as our primary prompt-based methodology and follow the framework presented in HiDe-Prompt [49]. During training, HiDe-Prompt co-optimizes task-specific prompts $\bm{p}_{t}$ and model’s output layer parameters $\psi$ for each new task $t$ using the WTP objective. These prompts are stored within a prompt pool $\mathbf{P}=\{\bm{p}_{1},...,\bm{p}_{T}\}$. At test time, a separate lightweight auxiliary output layer $\hat{h}_{\omega}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{T}$, trained with the TII objective, takes the uninstructed representation $f_{\theta}(x)$ of a new data point $\bm{x}$ as input to infer the task identity, guiding the selection of the most suitable prompt $\bm{p}_{k}$ from the prompt pool $\mathbf{P}$. Subsequently, the final prediction is given as $\hat{y}=h_{\psi}(f_{\theta}({\bm{x}},\bm{p}_{k}))$. For further details, please refer to Appendix D.
@@
where $G({\bm{h}}):=\mathrm{softmax}(s_{1}({\bm{h}}),\ldots,s_{N}(h))$. Building on this concept, works by [10, 41] established the MoE layer as a fundamental building block to scale up model capacity efficiently. Please refer to Appendix C for a comprehensive discussion of related works.
@@ 3.1 Mixture of Experts Meets Attention
Figure 1: An illustrative depiction of the relationship between self-attention and MoE. Each output vector of a head in the MSA layer can be viewed as the output of a MoE model. These MoE models share the same set of experts encoded in the value matrix. Each entry in the attention matrix corresponds to a score function within this architecture.
@@
for $i\in[N]$. Expanding on equation (8), we can discern that each attention head within the MSA layer implicitly embodies a special mixture of experts architecture. This architecture encompasses $N$ MoE models, each featuring its own quadratic gating function $G_{i}$. However, instead of employing $N^{2}$ separate expert networks for each model, this architecture utilizes $N$ shared linear expert networks $f_{j}$ for $j\in[N]$, significantly reducing the number of parameters. Notably, each expert network and its corresponding gating function process the entire input sequence directly, rather than individual embedding ${\bm{x}}_{i}$ as in traditional MoE layers [41]. This connection between self-attention and mixture of experts is depicted in Figure 1. In the subsequent section, we explore how prompt-based techniques can be viewed through this lens.
@@ 3.2 Prefix Tuning via the Perspective of Mixture of Experts
Figure 2: Left: An illustrative depiction of prefix tuning as the introduction of new experts into pre-trained MoE models. Right: Visualization of NoRGa implementation, integrating non-linear activation and residual connections into the prefix tuning attention matrix.
Building on the connection between self-attention and mixture of experts, we propose that applying prefix tuning can be interpreted as the introduction of new experts to customize the pre-trained model for a specific task, as illustrated in Figure 2. Specifically, similar to Section 3.1, we consider the $l$-th head within the MSA layer. We denote $\bm{p}^{K}=\left[\bm{p}_{1}^{K},\dots,\bm{p}_{L}^{K}\right]^{\top}\in\mathbb{R}^{L\times d}$, $\bm{p}^{V}=\left[\bm{p}_{1}^{V},\dots,\bm{p}_{L}^{V}\right]^{\top}\in\mathbb{R}^{L\times d}$, where $L=\frac{L_{p}}{2}$. We define new prefix experts $f_{N+j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$ along with their corresponding new score functions $s_{i,N+j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}$ as follows:
@@
We argue that our introduction of a novel connection between self-attention, prefix tuning, and MoE offers a fresh perspective on the design of previous prompt-based continual learning methods. In the context of continual learning, the pre-trained experts serve as a knowledge base, while prefix tuning augments it with task-specific knowledge encoded in new experts. Moreover, we draw a parallel between the pre-trained experts and the G(eneral)-Prompt utilized in DualPrompt, which captures task-agnostic information [53]. Both are shared across tasks, making them useful for prediction, especially when task identification is incorrect. Notably, the new experts achieve their efficiency through simple linear gating functions and independence from the input, unlike the pre-trained experts. For simplicity, we call the MoE model (11) as linear gating prefix MoE.
@@ 4 Non-linear Residual Gate Meets Prefix Tuning
As discussed earlier, prefix tuning introduces additional experts within the MoE framework, resulting in the linear gating prefix MoE model. However, as outlined in Appendix A, this approach suffers from suboptimal sample efficiency for parameter estimation. To overcome this and enhance overall CIL performance, we propose an innovative approach that significantly improves sample efficiency while promoting WTP performance in Section 4.1 and provide theoretical explanations in Section 4.2.
@@ 4.1 NoRGa: Non-linear Residual Gate
where $\alpha,\tau\in\mathbb{R}$ are learnable scalar factors, and $\sigma$ is a non-linear activation function. The new score function in equation (12) consists of a linear and a non-linear component. We call the new prefix MoE model with score functions (12) as non-linear residual gating prefix MoE.
The use of a non-linear activation function here is motivated by the algebraic independence condition in Definition 4.2 to theoretically guarantee the optimal sample efficiency of expert and parameter estimations (cf. Theorem 4.3). It’s worth noting that removing the linear component $s_{i,N+j}({\bm{X}})$ in the score function (12) could potentially lead to the vanishing gradient problem during training. To mitigate this challenge, we incorporate a residual connection [14] into the formulation. Our modification introduces minimal additional parameters ($\alpha$ and $\tau$) compared to the original score function, ensuring parameter efficiency. This is particularly crucial in continual learning scenarios where the number of parameters grows with each new task. For implementation, we define $H_{l}=W_{l}^{Q}{W_{l}^{K}}^{\top}$. From equation (5), the attention matrix of the $l$-th head can then be written as:
$A_{l}=\frac{{\bm{X}}^{Q}H_{l}[{\bm{p}^{K}}^{\top},\ {{\bm{X}}^{K}}^{\top}]}{\sqrt{d_{v}}}=\frac{[{\bm{X}}^{Q}H_{l}{\bm{p}^{K}}^{\top},\ {\bm{X}}^{Q}H_{l}{{\bm{X}}^{K}}^{\top}]}{\sqrt{d_{v}}}=[A_{l}^{\mathrm{prompt}},\ A_{l}^{\mathrm{pretrain}}].$ (13)
Here, $A_{l}^{\mathrm{prompt}}$ denotes the attention score matrix for the prompts, and $A_{l}^{\mathrm{pretrain}}$ represents the attention score matrix for the pre-trained experts. To implement NoRGa, we can directly modify the final attention matrix as follows:
$\hat{A}_{l}$ $=[\hat{A}_{l}^{\mathrm{prompt}},A_{l}^{\mathrm{pretrain}}],$ (14)
$\hat{A}_{l}^{\mathrm{prompt}}$ $=A_{l}^{\mathrm{prompt}}+\alpha\cdot\sigma(\tau\cdot A_{l}^{\mathrm{prompt}}).$ (15)
The implementation of NoRGa is illustrated in Figure 2. Despite its simplicity, our modification can significantly enhance sample efficiency and promote more reasonable parameter estimation, as demonstrated in our theoretical analysis in Section 4.2. Within the HiDe-Prompt framework, task-specific prompt parameters are trained using the WTP objective for each new task. Consequently, our modification leads to better parameter estimation, which directly contributes to improved WTP performance, ultimately improving overall continual learning efficacy. Importantly, NoRGa maintains the same parameter count as HiDe-Prompt, which is crucial in CL because of the memory constraint. Here, we evaluated $\sigma$ with $\tanh$, $\mathrm{sigmoid}$, and $\mathrm{GELU}$, finding $\tanh$ to perform well in most cases.
@@ 5 Experiments
Datasets. We evaluate various continual learning methods on widely used CIL benchmarks, including Split CIFAR-100 [23] and Split ImageNet-R [23], consistent with prior work [49]. We further explore the model’s performance on fine-grained classification tasks with Split CUB-200 [48] and large inter-task differences with 5-Datasets [9]. Please refer to Appendix E for more details.
Evaluation Metrics. We utilize several established metrics described in [50]. These include: final average accuracy (FA), which represents the average accuracy after the final task; cumulative average accuracy (CA), which refers to the historical average accuracy; and average forgetting measure (FM). We give more emphasis to FA and CA due to their comprehensiveness, as noted in [42].
Baselines. We compare our approach with several representative prompt-based approaches including L2P [54], DualPrompt [53], CODA-Prompt [42], S-Prompt [52], and HiDe-Prompt [49]. Additionally, we evaluate against state-of-the-art pre-trained model-based continual learning methods, including ADAM [60] and RanPAC [29]. We further extend our evaluation by applying HiDe-Prompt with parameter-efficient fine-tuning techniques like LoRA [17] and Adapters [16]. In line with [49], we utilize the checkpoints of ViT that use supervised pre-training of Imagenet-21K (denoted as Sup-21K), and some self-supervised pre-training such as iBOT-21K, iBOT-1K [62], DINO-1K [4], and MoCo-1K [7]. For implementation details, see Appendix E.
@@
Main Results. In Table 1, we evaluate several continual learning methods on Split CIFAR-100 and Split ImageNet-R using diverse pre-trained models. NoRGa achieves state-of-the-art FA and CA across all datasets and models, consistently outperforming HiDe-Prompt. On Sup-21K, NoRGa demonstrates impressive FA results on both CIFAR-100 and ImageNet-R. It also maintains the highest CA, with significant margins of 1.80% and 2.92% on CIFAR-100 and ImageNet-R, respectively, compared to HiDe-Prompt. These results highlight NoRGa’s strong ability to retain knowledge and exhibit minimal forgetting, as evidenced by the low FM values on both datasets. NoRGa also surpasses HiDe-Prompt on self-supervised models, with FA improvements up to 1.95% and 3.66%. We further investigate two scenarios: fine-grained classification tasks and large inter-task differences through experiments on Split CUB-200 and 5-Datasets, respectively, as shown in Table 2. NoRGa maintains its lead, achieving FA gaps of 4.34% and 2.46% on Split CUB-200, and the highest FA on 5-Datasets. While gains in some metrics may be modest, NoRGa consistently outperforms HiDe-Prompt in either FA or CA, underscoring its robustness. For example, on Split ImageNet-R with Sup-21K weights, the FA improvement is small (75.06% vs. 75.40%), but the CA gains are substantial (76.60% vs. 79.52%), demonstrating the method’s effectiveness and robustness.
Ablation Study. To assess the impact of non-linear activation functions on NoRGa’s performance, we evaluated the model’s behavior with different choices for the activation function $\sigma$, including $\tanh$, $\mathrm{sigmoid}$, and $\mathrm{GELU}$ in Table 3. The results show that NoRGa achieves state-of-the-art performance on both Split CIFAR-100 and Split CUB-200 datasets with all three activation functions. These findings suggest that NoRGa exhibits robustness to the choice of non-linear activation within a reasonable range. While all functions perform well, the $\tanh$ activation function demonstrates generally strong performance across scenarios. Further results are provided in the Appendix.
@@ 6 Conclusion
Supplement to “Mixture of Experts Meets Prompt-Based Continual Learning”
In this supplementary material, we first analyze the statistical suboptimality of the Linear Gating Prefix MoE Model (11) in Appendix A. Appendix B provides proofs for the theoretical results presented in Section 4.2. In Appendix C, we discuss related works on mixture of experts. Appendix D outlines the training algorithm for HiDe-Prompt while Appendix E presents the experimental setup and details. Further, Appendix F presents further experiments on the task-incremental learning setting to empirically demonstrate the benefits of using our proposed Non-linear Residual Gating Prefix MoE (12) over the Linear Gating Prefix MoE Model. Appendix G and Appendix H compare NoRGa with other parameter-efficient fine-tuning techniques and pre-trained model-based methods. In Appendix I, we present the efficiency tests, while Appendix J explores the impact of learnable $\alpha$ and $\tau$. Finally, Appendix K compares the training times between NoRGa and HiDe-Prompt.